A small stone of mass m, is thrown vertically upward with initial speed ,u. If the air resistance at speed v is mkv2, where k is a positive constant. Show that the stone returns to its starting point with speed given by v=u/((1+(ku2/g))1/2)
1
Expert's answer
2020-01-22T04:23:18-0500
"W=-mg"

"F_r=-mkv^2"

"ma=m\frac{dv}{dt}=-m(g+kv^2)"

"v\frac{dv}{dx}=-(g+kv^2)"

"-2\int_0^0dx=\int_u^v\frac{dv^2}{g+kv^2}"

"v=u\left(1+\frac{ku^2}{g}\right)^{-\frac{1}{2}}"